40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12534     Entry time: Wed Oct 5 19:43:13 2016
Author: gautam 
Type: Summary 
Category: General 
Subject: Vent review 

This elog is meant to review some of the important changes made during the vent this summer - please add to this if I've forgotten something important. I will be adding this to the wiki page for a more permanent record shortly.


Vent objectives:

  1. Clean ITMX, ITMY, ETMX, ETMY
  2. Replace ETMX suspension cage, replace Al wire standoffs with Ruby (sapphire?) standoffs.
  3. Shorten Y arm length by 20mm
  4. Replace 40mm aperture baffles in ETM chambers with 50mm black glass baffles

Optics, OSEM and suspension status:

ITMX & ITMY

  • ITMX and ITMY did not have any magnets broken off during the vent - all five OSEM coils for both were removed and the optic EQ stopped for F.C. cleaning.
  • Both HR and AR faces were F.Ced, ~20mm dia area cleaned.
  • The coils were re-inserted in an orientation as close to the original (as judged from photos), and the shadow sensor outputs were made as close to half their open values as possible, although in the process of aligning the arms, this may have changed
  • OSEM filter existense was checked (to be updated)
  • Shadow sensor open values were recorded (to be updated)
  • Checked that tables were level before closing up
  • The UL OSEM on ITMY was swapped for a short OSEM while investigating glitchy shadow sensor outputs. This made no difference. However, the original OSEM wasn't replaced. Short OSEM was used as we only had spare short OSEMs. Serial number (S/N 228) and open voltage value have been recorded, wiki page will be updated. Does this have something to do with the input matrix diagonalization weirdness we have been seeing recently?
  • ITMX seems to be prone to getting stuck recently, reason unknown although I did notice the LL OSEM was kind of close to the magnet while inserting (but this magnet is not the one getting stuck, as we can see this clearly on the camera - the prime suspect is UL I believe)
  • OL beam centering on in vacuum steering optics checked before closing up

ETMY

  • UL, UR and LR magents broke off at various points, and so have been reglued
  • No standoff replacement was done
  • Re-suspension was done using newly arrived SOS wire
  • Original OSEMs were inserted, orientations have changed somewhat from their previous configuration as we did considerable experimentation with the B-R peak minimization for this optic
  • OSEM filter status, shadow sensor open voltage values to be updated.
  • New wire suspension clamp made at machine shop is used, 5 in lb of torque used to tighten the clamp
  • HR face cleaned with F.C.
  • Optic + suspension towers air baked (separately) at 34C for curing of EP30
  • Checked that tables were level before closing up
  • 40mm O.D. black glass baffle replaced with 50mm O.D. baffle.
  • Suspension cage was moved towards ITMY by 19mm (measured using a metal spacer) by sliding along stop marking the position of the tower.

ETMX

  • Al wire standoffs <--> Ruby wire standoffs (this has changed the pitch frequency)
  • All magnets were knocked off at some point, but were successfully reglued
  • New SOS tower, new SOS wire, new wire clamp used
  • OSEM filter status, shadow sensor open voltage values to be updated.
  • OSEM orientation is close to horizontal for all 5 OSEMs
  • Table leveling was checked before closing up.
  • 40mm O.D. black glass baffle replaced with 50mm O.D. baffle.\

PRM

  • Some issues with the OSEMs were noticed, and were traced down to the Al foil caps covering the back of the (short) OSEMs, which are there to minimize the scattererd 1064nm light interfering with the shadow sensor, shorting one of the OSEMs
  • To mitigate this, all Al foil caps now have a thin piece of Kapton between foil and electrical contacts on rear of OSEM
  • No OSEMs were removed from the suspension cage during this process, we tried to be as gentle as possible and don't believe the shadow sensor values changed during this work, suggesting we didn't disturb the coils (PRM wasn't EQ stopped either)

SRM

  • The optic itself wasn't directly touched during the vent - but was EQ stopped as work was being done on ITMY
  • It initially was NOT EQ stopped, and the shift in table level caused by moving ITMY cage to the edge of the table for F.C. cleaning caused the optic to naturally drift onto the EQ stops, leading to some confusion as to what happened to the shadow sensor outputs
  • The problem was diagnosed and restoring ITMY to its original position made the OSEM signals come back to normal.

SR3

  • Was cleaned by drag wiping both front and back faces

SR2/PR2/PR3/BS/OMs

  • These optics were NOT intentionally touched during this vent
  • The alignment on the OMs was not checked before close-up
 

Other checks/changes

  • OL beams were checked on in-vacuum input and output steering mirrors to make sure none were close to clipping
  • Insides of viewport windows were checked for general cleanliness, given that we have found the outside of some of these to be rather dirty. Insides of viewports checked were deemed clean enough.
  • Steve has installed a new vacuum guage to provide a more realiable pressure readout. 
  • We forgot to investigate the weird behaviour of the AS beam that Yutaro and Koji identified in November. In any case, looks like the clipping of the AS beam is worse now. We will have to try and fix this using the PZT mounted OMs, and if not, we may have to consider venting again

Summary of characterization tasks to be done:

  1. Mode matching into the Y arm cavity given the arm length change
  2. HOM content in transmitted IR light from Y arm given the arm length change (Finesse models suggest that the 2f second order HOM resonance may have moved closer to the 00 resonance)
  3. Arm loss measurement
  4. Suspension diagonalization
  5. Check the Qs of the optics eigenmodes - should indicate if any of our magnets, reglued or otherwise, are a little loose
ELOG V3.1.3-